OpenStudy (anonymous):

Find the volume of the solid bounded by the cylinder y^2 + z^2 = 4 and the planes x = 2y, x=0, z=0 in the first octant

OpenStudy (abb0t):

\[\int\limits_{0}^{2} \int\limits_{0}^{2y}\sqrt{4-y^2}dV = \int\limits_{0}^{2} \int\limits_{0}^{2y}\sqrt{4-y^2}dxdy \]

OpenStudy (abb0t):

The rest is simple integration using u-substitution.

OpenStudy (anonymous):

Thanks I could set it up as a triple integral, but didn't understand a double integral